Aggregator
Autolog save
QuickPic - MBC3A11H mismatch - setpoint taken and cycled
ATLis Comments Daily Digest
Log Number: 4367989Logbook(s): Tags: Body:
*/
ATLis Comments Posted 18:00 Tuesday April 29 through 18:00 Wednesday April 30
AES DC power Systems
Replace Utility Chassis in NA01B10 Trim Rack
2025-04-30 16:04 - Eric Diggs
TAGGERB has been enabled in alarm handler
"Hall B: Tagger Turning On"
QuickPic - Fifth pass horizontal extraction
Injector Alarm Limits Set
Injector Alarm Limits Set
QuickPic - MQABE04 mismatch - setpoint taken and cycled
ioctsbdiag restored
Log Number: 4367967Logbook(s): Tags: Body: restorelib 2.6
Target restore date was 2025-04-30 16:38:40
The output from the myRestore execution was as follows:
ioctsbdiag
vac@2025-04-30 16:38
11(11) restored
systemEnum@2025-04-30 16:38
2(2) restored
Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -i ioctsbdiag (success)
PVs modified: 0
PVs modified: 0
PVs modified: 0
PVs modified: 0
PVs modified: 0
PVs modified: 0
Procedure return status: Success
Operator SIGN IN
Operator SIGN IN
6E02 viewer (Hall A,B,C,D)
Phaser Results
COMMENT ON Replace Utility Chassis in NA01B10 Trim Rack
Log Number: 4367945Logbook(s): Body:
We replace the Utility Chassis in Trim Rack NA01B10 and had Ops cycle all the magnets. We will leave the ATLis and PTW open from a week, to see if the Trim Rack trip off again. If it does trip off again, we may need facilities to look at the breaker panel for this Trim Rack.
dogCalc V0-13 download
2L22 HV Drop!!! *o*
Vacuum Group Summary 2025-04-30
Fopt - Prior to Pass change
Log Number: 4367929Logbook(s): Tags: Body: Fopt 7.9 completed successfully. The runtime settings include a sample count of 15, a minimum current of 4, and an operating regime of Standard with Excitation in 1S Region. The corrector and RF kicks applied were 1.50003 and 1.50003 volts respectively. The max. amplitude setting was 1 millimeters. The files for this run were saved in the directory /cs/opshome/hlapps/optics/diforb/25apr301541. The X and Y stability of the machine during the run was 0.0502639 and 0.0320722 (mm) respectively.